GlueX TOF Meeting, September 12, 2013

From GlueXWiki
Jump to: navigation, search

GlueX Time-of-Flight Meeting
Thursday, September 12, 2013
10:00 am EDT
JLab: CEBAF Center, F326/327


  1. Announcements
  2. Review of meeting on August 29
  3. October Collaboration Meeting
  4. Construction news [1][2]
  5. Installation plan



  1. ESNet: 8542553

Desktop Sharing

  1. Go to
  2. In the "join a meeting" box enter the code: 1833622
  3. Fill in the participant registration form.


To connect by telephone:

  1. dial:
  2. enter access code followed by the # sign: 1833622#


Talks can be deposited in the directory /group/halld/www/halldweb/html/talks/2013-3Q on the JLab CUE. This directory is accessible from the web at .



  • FSU: Hussein Al Goul, Paul Eugenio, Sasha Ostrovidov, Aristedis Tsaris
  • JLab: Chuck Hutton, Mark Ito (chair), Beni Zihlmann

Review of Minutes from the August 29 Meeting

We reviewed the minutes from last time.

Chuck updated us on status of pieces for the support structure to be machined at FSU.

  • Some of the drawings for have not been reviewed yet, but will be soon. The important ones have been signed; this is not holding us up.
  • The material for the aluminum collars have been ordered and should be on the way to FSU.
  • Randall Smith has contacted Chuck about the length of the aluminum "cheese pieces", those that hold the light guides near the PMT. The FSU shop cannot make them 40" long as specified. Chuck will review the design and try two 20" pieces instead. When he is done he will back to Randall to specify the order for raw materials.

Collaboration Meeting

Sasha will give the talk at the collaboration meeting on TOF construction status.

Construction News

Paul gave an update on the status of construction.

  • Assembly is complete on the bulk of the standard modules.
    • There are four remain standard modules that need PMTs and wrapping.
    • There are 12 modules that only lack the final Tedlar wrapping.
    • Four modules need re-working of the joint between the light guide and PMT.
      • This issue is solved now. The latest modules have clean looking glue joints.
    • By the end of the month all of standard modules should be done.
  • The next major tasks are construction of the eight half-width and eight half-length modules.
  • In October, FSU will turn to packing the modules and the details of transport to JLab. They will likely drive them up in two batches, one layer at a time.
  • Chuck will check with Tim on storage space at JLab.
  • FSU students are polishing the spare light guides that were machined, in case they are needed.
  • Next week Maurizio will lead a search for extremely straight lumber to be used in the shipping crates.

Module Testing

Sasha gave the report.

  • Seven batches of modules have been tested. Testing of batch 8 is in progress. Recall that each batch consists of six standard modules.
  • He is now using the TDC lent by JLab. The FSU modules has been sent back to CAEN for repair.
  • The problem reported last time with trigger jitter has come and gone several times during the past two weeks. We have arranged to ship another discriminator to FSU from JLab. It will be tried to see if that changes behavior. The discriminator will be shipped as soon as Chris Stanislav returns from vacation.
  • Sasha showed an updated plot of nominal gain (at standard HV, from the Hamamatsu spec sheet) vs. the HV arrived at after the gain balancing procedure. All tubes in the first seven batches are represented. As expected, tubes with low nominal gain require a higher voltage operation point to be in balance with other tubes.
  • He also showed a plot of nominal dark current (again from the manufacturer's spec sheet) vs. singles counting rate for each PMT. Here he correlation is weak if it exists at all.
  • For the 42 modules test thus far:
    • The speed-of-light measurements are consistent among the modules with an average of 15.2 cm/ns.
    • Mean time resolution is also consistent in the low to mid 90's, measured in picoseconds. The one batch that was measured when the trigger time jitter problem was present is in the high 90's. The measurements for that batch may have to be repeated.
    • Position resolution results have not been tabulated yet.
  • There is still development to do on the test procedure for the single-ended counters (half-length).